OpenStudy (anonymous):

Hunus: A sign at Pablo's Pizza says that the max number of people it can serve is 225 customers. The restaurant has 3 large tables which seat 12 people at each and x tables which seat 6 people each. What is the inequality that can be used to determine the number of tables for 6 in the restaurant?

OpenStudy (anonymous):

a. 6x+12≥25 b. 225+6x<36 c. 36+6x≥225 d. 6x-24<225 e. 6x+36≤225

OpenStudy (hunus):

\[(3)(12) + 6x \le 225\] \[36 +6x \le 225\] \[6x \le 189\] \[x \le \frac{189}{6}\] \[x \le \frac{63}{2}\]

OpenStudy (anonymous):

you are a lifesaver.

OpenStudy (hunus):

:)

OpenStudy (anonymous):

I have like 4 questions more. Can you help me with them? If you're too busy, it's fine :)

OpenStudy (hunus):

Sure

OpenStudy (anonymous):

Thankyou! I appreciate it so muchhh

OpenStudy (hunus):

It's no problem :)